OpenStudy (anonymous):

Find the greatest common factor of these two expressions. 18x^5y^2 and 27x ^5y^8w^3

OpenStudy (michele_laino):

we can rewrite your monomials as follows: \[\Large \begin{gathered} 18{x^5}{y^2} = 2 \times {3^2}{x^5}{y^2} \hfill \\ \\ 27{x^5}{y^8}{w^3} = {3^3}{x^5}{y^8}{w^3} \hfill \\ \end{gathered} \]

OpenStudy (michele_laino):

now, you have to pick all common factors with the lowest exponent

OpenStudy (michele_laino):

for example: is 2 a common factor?

OpenStudy (anonymous):

the correct answer is 9x^5y^2

OpenStudy (michele_laino):

that's right!